please help mark had $20 before he began earning the same amount each week at his new job he used half of his first week's pay to pay back a loan. he spent $15 at the movies and $12 on a book and got $25 for his birthday after his second paycheck he had$20 less than two weeks pay combined how much is his paycheck?
Found 2 solutions by sofiyac, edjones:
Answer by sofiyac(983)

mark had $20 before he began earning the same amount each week at his new job he used half of his first week's pay to pay back a loan. he spent $15 at the movies and $12 on a book and got $25 for his birthday after his second paycheck he had$20 less than two weeks pay combined how much is his paycheck?
If he earns x amount each week for 2 weeks then the total money he has would be 2x+20. Now we can subtract anything he spent. Half of first week's pay would be x/2 so we have
2x+20-x/2-15-12+25 so this what he has after second paychecks and problems says that it was $20 less than two weeks pay so 2x-20. Now we can set them equal to each other and solve for x.
2x+20-x/2-15-12+25=2x-20
-x/2+18=-20
-x/2=-38
x=76 So he makes $76 per week.

Answer by edjones(8007)

Let x = amount of each paycheck.
20+2x - x/2 -15-12+25=2x-20
-x/2=-38
x=$76.
.
Ed